# Jailbreak Attack Initializations as Extractors of Compliance Directions

### v2

@@ Abstract
Safety-aligned LLMs respond to prompts with either compliance or refusal, each corresponding to distinct directions in the model’s activation space. Recent works show that initializing attacks via self-transfer from other prompts significantly enhances their performance. However, the underlying mechanisms of these initializations remain unclear, and attacks utilize arbitrary or hand-picked initializations. This work presents that each gradient-based jailbreak attack and subsequent initialization gradually converge to a single compliance direction that suppresses refusal, thereby enabling an efficient transition from refusal to compliance. Based on this insight, we propose $CRI$, an initialization framework that aims to project unseen prompts further along compliance directions. We demonstrate our approach on multiple attacks, models, and datasets, achieving an increased attack success rate ($ASR$) and reduced computational overhead, highlighting the fragility of safety-aligned LLMs. A reference implementation is available at [Uncaptioned image].
@@
Amit LeVi*,1, Rom Himelstein*,2, Yaniv Nemcovsky*,1, Avi Mendelson1, Chaim Baskin3 1Department of Computer Science, Technion - Israel Institute of Technology, 2Department of Data and Decision Science, Technion - Israel Institute of Technology, 3School of Electrical and Computer Engineering Engineering, Ben-Gurion University of the Negev Correspondence: amitlevi@campus.technion.ac.il, romh@campus.technion.ac.il.
**footnotetext: Equal contribution.
@@ 1 Introduction
A common strategy to enhance the safety of LLMs is safety-alignment, which involves training models to generate outputs that adhere to desired safety and ethical standards Shen et al. (2023); Wu et al. (2024b); Wang et al. (2023); Lee et al. (2023). This method distinguishes between harmless and harmful prompts to determine whether they should be complied with or refused Glaese et al. (2022); Wang and Isola (2020). Thereby, effectively segmenting the input space into Compliance and Refusal subspaces Yu et al. , where previous works have shown that each subspace correlates to distinct directions within the LLM’s internal activation space Arditi et al. (2024); Wollschläger et al. (2025). Inadvertently, this fuels jailbreak attacks that manipulate harmful prompts to elicit models into compliance, contrary to their safety guidelines Marshall et al. (2024); Yu et al. (2024b); Chao et al. (2023); Deng et al. (2023).
@@
Gradient-based jailbreak attacks are a well-known category that utilizes gradient descent optimization schemes via models’ backpropagation Zou et al. (2023); Liu et al. (2023); Zheng et al. (2024); Schwinn et al. (2024). Such attacks follow a white-box setting, where attackers can access models’ gradients, and have been shown to transfer across different models Wu et al. (2025a); Zou et al. (2023); Liu et al. (2023); ATTACKS (2025). This transferability property reflects the shared vulnerabilities of LLMs Zhao et al. (2025b), and extends the attacks’ relevance to black-box settings, where surrogate models are used in the optimization Li et al. (2024c); Zhang et al. (2025).
@@
This work introduces Compliance Refusal Initialization ($CRI$), a novel initialization framework for gradient-based jailbreak attacks. Our approach considers the activation-space refusal direction defined by Arditi et al. (2024) as the average difference between refusal and compliance prompts-induced LLM activations. Previous works showed that while refusal is governed by a single direction, there are multiple opposing compliance directions utilized by jailbreak attacks Arditi et al. (2024); Ball et al. (2024); He et al. (2024); Zhou et al. (2024b). Nonetheless, the optimization of given jailbreak attacks often exhibits similar compliance directions over different prompts Ball et al. (2024). $CRI$ then leverages pre-trained jailbreak attacks to project unseen prompts along these compliance directions, and towards the compliance subspace. In Figure 1, we illustrate the initializations’ effect using a compliance-refusal SVM classifier, where $CRI$ projects prompts toward the decision boundary. We provide the detailed configurations and additional analysis in Section A.1. Below, we outline our main contributions.
@@
  • We propose the $CRI$ framework, which pre-computes self-transfer initializations and utilizes the compliance directions’ similarity to identify and utilize effective initializations over given prompts.
@@
The rest of the paper is organized as follows: Section 2 discusses the attack setting and related works, Section 3 describes our proposed method, Section 4 provides our experimental results, Section 5 concludes the paper, and Section 6 discusses the limitations of our work.
@@ 2.1 Background
The predefined set of jailbreak transformations $JT$ can be considered to limit the attack to transformations that preserve the intention in the input prompt $x$. A common practice in text-based adversarial attacks is to consider transformations that add a textual suffix and or a prefix to the input with up to a given length $L$ Zou et al. (2023); Liu et al. (2023); Yu et al. ; IGCG et al. (2024); Guo et al. (2024). Embedding-based attacks utilize similar suffix transformations but extend the scope of possible suffixes to any embedding vectors rather than those that align with textual tokens Schwinn et al. (2024). Accordingly, such attacks refer to the token embeddings of LLMs as the input representations [^1] This can be extended to include attacks on multilayer representations, as in Schwinn et al. (2024).. Depending on the transformation type, the attack then optimizes and transforms the discrete text or continuous embeddings. Formally:
@@
Figure 2: Comparison of directions’ cosine-similarity during $GCG$’s optimization process on the $HarmBench$ dataset over the Llama-2 model. We compare the refusal with attacks and self-transfer initializations (left), and present the similarity matrices of attacks (center) and initializations (right).
@@ 2.2 Related Work
Self-transfer-based initialization was first introduced by Andriushchenko et al. (2024) as a strategy for enhancing random-search jailbreak attacks, demonstrating improvements in $ASR$ and reducing computational overhead. Building on this idea, IGCG et al. (2024) identified a category of malicious prompts that provided more effective initializations. Other studies, such as Schwinn et al. (2024) and Liu et al. (2023), employed hand-crafted jailbreaks as initializations and demonstrated further enhanced effects. Our work extends these approaches by building on theoretical insights that enable the automatic recognition of effective initializations over given prompts.
@@ 3.1 Motivation
In this section, we discuss the motivation of the $CRI$ framework. We first consider that attacks project prompts in opposing directions of refusal and toward similar compliance directions Ball et al. (2024). We expand on this phenomenon by considering the gradual optimization of given gradient-based attacks over different prompts. In Figure 2, we present that such attacks, and subsequent self-transfer initializations, show highly similar directions during their optimization process, while shifting further away from refusal. This entails that gradient-based attacks gradually transfer from refusal to compliance by following a single direction, suggesting that effective initializations project prompts further along this direction. We denote this direction as the compliance direction, which is defined independently for each attack and prompts’ distribution. As the attack optimization aims to minimize the attack criterion, it indicates the progression along the compliance direction and thereby, the proximity to the compliance subspace. Therefore, we consider the metric of loss-in-the-first-step ($LFS$), which measures the attack criterion values when utilizing a given initialization over a given prompt. A lower $LFS$ then indicates that the initialization projects said prompt further along the compliance direction, which aids in guiding subsequent optimization and reduces the distance to a successful attack. $CRI$ then utilizes $LFS$ to select an effective initialization from a pre-computed set.
@@ 3.2 CRI Framework
$\mathcal{T}_{N\text{-}CRI}$ =$=\{A(x,t)\}_{(x,t)\in S_{FT}}$ (13)
$\mathcal{T}_{1\text{-}CRI}$ =$=\{A^{U}(S_{FT})\}$ (14)
@@
The single attack in $\mathcal{T}_{1\text{-}CRI}$ considers multiple prompts and is optimized for generalization to unknown prompts. However, it does not consider the $LFS$ metric and the corresponding attunement to the initialized prompt. In contrast, $\mathcal{T}_{N\text{-}CRI}$ optimizes each attack to achieve the minimum over a different prompt in Equation 12, achieving lower loss over the fine-tuning set and providing various initial attacks, which may be relevant to different prompts in deployment. However, these two approaches present corresponding disadvantages. $\mathcal{T}_{1\text{-}CRI}$ only contains a single attack and cannot address different behaviors in deployment. Contrarily, $\mathcal{T}_{N\text{-}CRI}$ separately optimizes each attack on a single prompt and does not consider its generalization to unknown prompts. Therefore, we extend $\mathcal{T}_{K\text{-}CRI}$ to consider a combination of the approaches via prompt clustering. We then use a universal attack on each cluster and produce multiple attacks with enhanced generalization properties.
@@
We now consider a combination of the approaches presented in $\mathcal{T}_{N\text{-}CRI},\mathcal{T}_{1\text{-}CRI}$ by grouping prompts towards applying $A^{U}$. Firstly, we note that the optimization in Equation 12 only considers the $LFS$ metric. However, $CRI$ benefits the entire attack optimization, especially if the initial starting point does not significantly change during the optimization. Hereby, initial attacks that generalize better to unseen prompts may remain partly unchanged and present a better initialization. Therefore, we utilize a fixed number of prompts to produce each attack in a given $CRI$ set, considering attacks with comparable generalization properties in deployment. In addition, we do not utilize a prompt for optimizing multiple attacks, as it may result in over-fitting to similar prompts. Each attack in $\mathcal{T}_{K\text{-}CRI}$ is then trained over $\left\lfloor\nicefrac{{N}}{{K}}\right\rfloor$ distinct prompts, and we correspondingly select $K$ prompt clusters. We aim to cluster similar prompts to initialize other similar prompts effectively. Hereby, we utilize a sentence encoder $EN$ Warner et al. (2024) and cluster prompt embeddings with minimal Euclidean distance. For clustering, we utilize the constrained k-means algorithm suggested by Bradley et al. (2000), where a single attack is then optimized over each cluster. The full algorithm is presented in Algorithm 1, and a visualization of the prompt clusters in Figure 8.
@@ 4 Experiments
This section presents a comprehensive empirical evaluation of $CRI$ compared to baseline initializations. We first present the experimental setting in Section 4.1, and continue to discuss the results over the $HarmBench$ dataset in Section 4.2. Additionally, we present ablation studies of our method in Section C.5, and extend the evaluations with additional attacks, models, and datasets in Section C.4. Our evaluation seeks to addresss three key research questions:
@@ 4.1 Experimental Setting
We present our experiments on the "standard behaviors" category of the $HarmBench$ dataset Mazeika et al. (2024), which is comprised of prompt-target pairs $(x,t)$. We define three disjoint subsets and utilize them in all the presented settings. $CRI$ utilizes $2$ subsets of $25$ samples each: a fine-tuning set for optimizing $CRI$, and a validation set for evaluating the universal attacks. The final subset is the $50$-sample test set over which we report the results. Additionally, we present results over the $AdvBench$ dataset Zou et al. (2023) in Sections C.4.3 and C.4.1.
Figure 3: Correlation of $LFS$ and steps-to-success on the $HarmBench$ dataset for the $GCG$ attack over the Llama-2 model across initializations. We present the linear regression over all samples (Pearson $r=0.46$, $p=7\times 10^{-8}$).
@@
Figure 4: Comparison of $ASR$ and number of steps on the $HarmBench$ dataset for the $GCG$ (top) and Embedding attacks (bottom) over the Llama-2 (left) and Vicuna (right) models.
@@
We evaluate our results over the open-source Vicuna-7B-v1.3 Chiang et al. (2023b), and Llama-2-7B-Chat-HF Touvron et al. (2023b) LLMs, which we accordingly denote as Vicuna, and Llama-2. We present these models to compare with the default setting presented by $GCG$, and provide results over additional models in Section C.1.
@@
We present the $GCG$ Zou et al. (2023) and the Embedding attacks Schwinn et al. (2024), where we utilized the same parameter configuration in all the presented settings, including the initializations’ training. The parameter configuration is then the default parameters suggested by the authors, except for the batch-size $B$ parameters of $GCG$. Thereby, $GCG$ utilizes $500$ optimization steps and the $JT$ set in Equation 7 with $L=20$ tokens. Similarly, the Embedding attack utilizes $100$ optimization steps and the $JT$ set in Equation 9 with $L=20$ tokens. To present a more robust setting, we limit the computational resources of the $GCG$ attack and take the batch-size parameter of $GCG$ to be $B=16$ instead of the default $512$ value. This entails that only $B=16$ tokens are sampled as possible replacements in each iteration of $GCG$, emphasizing the effect of the compared initializations. Additionally, we present results with increased batch size in Section C.5.1.
@@
We compare the $K-CRI$ initializations over $K=1,5,25$ to existing approaches, i.e., "Standard" which utilizes repeated ’!!!’ tokens, "Random" which utilizes uniform random tokens, and "Baseline" which utilizes randomly selected self-transfer initializations similarly to Andriushchenko et al. (2024); Jia et al. (2024). We then refer to "Standard" and "Random" initializations as uninformative and to others as informative. For each attack, $CRI$ pre-trains the transformations on the designated data subsets and utilizes $LFS$ to select an initialization for a given prompt. The "Baseline" approach then utilizes the same initializations set as $25\text{-}CRI$, but selects each initialization uniformly.
@@
In each attack’s evaluation, we follow the default evaluation framework suggested by the authors to compute the $ASR$. Thereby, $GCG$ evaluates success as not generating refusal-list keywords, where we provide the list in Section C.1.4. In contrast, the Embedding attack requires generating the exact target. We additionally present an evaluation of the resulting attacks via a GPT-4 judge in Section C.1.4. In addition to the $ASR$ and $LFS$, we define two metrics for estimating the computational overhead of attacks: $MSS$—Median Steps to Success, $ASS$—Average Steps to Success. Hereby, a lower $MSS$/$ASS$ of initializations indicates their corresponding decrease in computational overhead.
@@
In all the presented settings, the computation was executed on Intel(R) Xeon(R) CPU and NVIDIA L40S GPU.

### v3

@@ Abstract
Safety-aligned LLMs respond to prompts with either compliance or refusal, each corresponding to distinct directions in the model’s activation space. Recent studies have shown that initializing attacks via self-transfer from other prompts significantly enhances their performance. However, the underlying mechanisms of these initializations remain unclear, and attacks utilize arbitrary or hand-picked initializations. This work presents that each gradient-based jailbreak attack and subsequent initialization gradually converge to a single compliance direction that suppresses refusal, thereby enabling an efficient transition from refusal to compliance. Based on this insight, we propose $CRI$, an initialization framework that aims to project unseen prompts further along compliance directions. We demonstrate our approach on multiple attacks, models, and datasets, achieving an increased attack success rate ($ASR$) and reduced computational overhead, highlighting the fragility of safety-aligned LLMs. A reference implementation is available at [Uncaptioned image].
@@
Amit LeVithanks: Equal contribution.,1, Rom Himelstein11footnotemark: 1,2, Yaniv Nemcovsky11footnotemark: 1,1, Avi Mendelson1, Chaim Baskin3 1Department of Computer Science, Technion - Israel Institute of Technology, 2Department of Data and Decision Science, Technion - Israel Institute of Technology, 3School of Electrical and Computer Engineering, Ben-Gurion University of the Negev Correspondence: amitlevi@campus.technion.ac.il, romh@campus.technion.ac.il.
@@ 1 Introduction
A common strategy to enhance the safety of LLMs is safety-alignment, which involves training models to generate outputs that adhere to desired safety and ethical standards Shen et al. (2023); Wu et al. (2024b); Wang et al. (2023); Lee et al. (2023). This method distinguishes between harmless and harmful prompts to determine whether they should be complied with or refused Glaese et al. (2022); Wang and Isola (2020). Thereby, effectively segmenting the input space into Compliance and Refusal subspaces Yu et al. (2024a), where previous works have shown that each subspace correlates to distinct directions within the LLM’s internal activation space Arditi et al. (2024); Wollschläger et al. (2025). Inadvertently, this fuels jailbreak attacks that manipulate harmful prompts to elicit models into compliance, contrary to their safety guidelines Marshall et al. (2024); Yu et al. (2024c); Chao et al. (2023); Deng et al. (2023).
@@
Gradient-based jailbreak attacks are a well-known category that utilizes gradient descent optimization schemes via models’ backpropagation Zou et al. (2023); Liu et al. (2023); Zheng et al. (2024); Schwinn et al. (2024). Such attacks follow a white-box setting, where attackers can access models’ gradients, and have been shown to transfer across different models Wu et al. (2025a); Zou et al. (2023); Liu et al. (2023); Lin et al. (2025). This transferability property reflects the shared vulnerabilities of LLMs Zhao et al. (2025b), and extends the attacks’ relevance to black-box settings, where surrogate models are used in the optimization Li et al. (2024c); Zhang et al. (2025).
@@
Figure 2: Illustration of the $K\text{-}CRI$ framework. First, NN harmful prompts are clustered into $K$ groups, and a candidate attack is trained for each cluster (left). When a new prompt arrives, CRICRI evaluates the LFS of the $K$ attacks and selects the one with the lowest value as initialization for a refined attack (right).
This work introduces Compliance Refusal Initialization ($CRI$), a novel initialization framework for gradient-based jailbreak attacks. Our approach considers the activation-space refusal direction defined by Arditi et al. (2024) as the average difference between refusal and compliance prompts-induced LLM activations. Previous works showed that while refusal is governed by a single direction, there are multiple opposing compliance directions utilized by jailbreak attacks Arditi et al. (2024); Ball et al. (2024); He et al. (2024); Zhou et al. (2024b). Nonetheless, the optimization of given jailbreak attacks often exhibits similar compliance directions over different prompts Ball et al. (2024). $CRI$ then leverages pre-trained jailbreak attacks to project unseen prompts along these compliance directions, and towards the compliance subspace. In Figure˜1, we illustrate the initializations’ effect using a compliance-refusal SVM classifier, where $CRI$ projects prompts toward the decision boundary. We provide the detailed configurations and additional analysis in Section˜A.1. Below, we outline our main contributions.
@@
  • We propose the $CRI$ framework, which pre-computes self-transfer initializations and utilizes the compliance directions’ similarity to identify and utilize effective initializations over given prompts. An illustration is presented in Figure˜2.
@@
The rest of the paper is organized as follows: Section˜2 discusses the attack setting and related works, Section˜3 describes our proposed method, Section˜4 provides our experimental results, Section˜5 concludes the paper, and Section˜6 discusses the limitations of our work.
@@ 2.1 Background
The predefined set of jailbreak transformations $JT$ can be considered to limit the attack to transformations that preserve the intention in the input prompt $x$. A common practice in text-based adversarial attacks is to consider transformations that add a textual suffix and or a prefix to the input with up to a given length $L$ Zou et al. (2023); Liu et al. (2023); Yu et al. (2024a); IGCG et al. (2024); Guo et al. (2024). Embedding-based attacks utilize similar suffix transformations but extend the scope of possible suffixes to any embedding vectors rather than those that align with textual tokens Schwinn et al. (2024). Accordingly, such attacks refer to the token embeddings of LLMs as the input representations [^1] This can be extended to include attacks on multilayer representations, as in Schwinn et al. (2024).. Depending on the transformation type, the attack then optimizes and transforms the discrete text or continuous embeddings. Formally:
@@
Figure 3: Comparison of directions’ cosine-similarity during GCGGCG’s optimization process on the $HarmBench$ dataset over the Llama-2 model. We compare the refusal with attacks and self-transfer initializations (left), and present the similarity matrices of attacks (center) and initializations (right).
@@ 2.2 Related Work
Self-transfer-based initialization was first introduced by Andriushchenko et al. (2024) as a strategy for enhancing random-search jailbreak attacks, demonstrating improvements in $ASR$ and reducing computational overhead. This initialization used an arbitrary previously successful jailbreak attack as the starting point for further attack optimization. Building on this idea, IGCG et al. (2024) identified a category of malicious prompts that provided more effective initializations when using self-transfer. An arbitrary initialization is then selected from this category. Other studies, such as Schwinn et al. (2024) and Liu et al. (2023), employed hand-crafted jailbreaks as initializations and demonstrated further enhanced effects. Our work extends these approaches by building on theoretical insights that enable the automatic selection of effective initializations over given prompts. We supplement these previous initialization methods, which were not designed to select self-transfer over individual prompts.
@@ 3.1 Motivation
In this section, we discuss the motivation of the $CRI$ framework. We first consider that attacks project prompts in opposing directions of refusal and toward similar compliance directions Ball et al. (2024). We expand on this phenomenon by considering the gradual optimization of given gradient-based attacks over different prompts. In Figure˜3, we present that such attacks, and subsequent self-transfer initializations, show highly similar directions during their optimization process, while shifting further away from refusal. This entails that gradient-based attacks gradually transfer from refusal to compliance by following a single direction, suggesting that effective initializations project prompts further along this direction. We denote this direction as the compliance direction, which is defined independently for each attack and prompts’ distribution. As the attack optimization aims to minimize the attack criterion, it indicates the progression along the compliance direction and thereby, the proximity to the compliance subspace. Therefore, we consider the metric of loss-in-the-first-step ($LFS$), which measures the attack criterion values when utilizing a given initialization over a given prompt. A lower $LFS$ then indicates that the initialization projects said prompt further along the compliance direction, which aids in guiding subsequent optimization and reduces the distance to a successful attack. $CRI$ then utilizes $LFS$ to select an effective initialization from a pre-computed set.
@@ 3.2 CRI Framework
$\mathcal{T}_{N\text{-}CRI}$ $=\{A(x,t)\}_{(x,t)\in S_{FT}}$ (13)
$\mathcal{T}_{1\text{-}CRI}$ $=\{A^{U}(S_{FT})\}$ (14)
@@
The single attack in $\mathcal{T}_{1\text{-}CRI}$ considers multiple prompts and is optimized for generalization to unknown prompts. However, it does not consider the $LFS$ metric and the corresponding attunement to the initialized prompt. In contrast, $\mathcal{T}_{N\text{-}CRI}$ optimizes each attack to achieve the minimum over a different prompt in Equation˜12, achieving lower loss over the fine-tuning set and providing various initial attacks, which may be relevant to different prompts in deployment. However, these two approaches present corresponding disadvantages. $\mathcal{T}_{1\text{-}CRI}$ only contains a single attack and cannot address different behaviors in deployment. Contrarily, $\mathcal{T}_{N\text{-}CRI}$ separately optimizes each attack on a single prompt and does not consider its generalization to unknown prompts. Therefore, we extend $\mathcal{T}_{K\text{-}CRI}$ to consider a combination of the approaches via prompt clustering. We then use a universal attack on each cluster and produce multiple attacks with enhanced generalization properties.
@@
We now consider a combination of the approaches presented in $\mathcal{T}_{N\text{-}CRI},\mathcal{T}_{1\text{-}CRI}$ by grouping prompts towards applying $A^{U}$. Firstly, we note that the optimization in Equation˜12 only considers the $LFS$ metric. However, $CRI$ benefits the entire attack optimization, especially if the initial starting point does not significantly change during the optimization. Hereby, initial attacks that generalize better to unseen prompts may remain partly unchanged and present a better initialization. Therefore, we utilize a fixed number of prompts to produce each attack in a given $CRI$ set, considering attacks with comparable generalization properties in deployment. In addition, we do not utilize a prompt for optimizing multiple attacks, as it may result in over-fitting to similar prompts. Each attack in $\mathcal{T}_{K\text{-}CRI}$ is then trained over $\left\lfloor\nicefrac{{N}}{{K}}\right\rfloor$ distinct prompts, and we correspondingly select $K$ prompt clusters. We aim to cluster similar prompts to initialize other similar prompts effectively. Hereby, we utilize a sentence encoder $EN$ Warner et al. (2024) and cluster prompt embeddings with minimal Euclidean distance. For clustering, we utilize the constrained k-means algorithm suggested by Bradley et al. (2000), where a single attack is then optimized over each cluster. The full algorithm is presented in Algorithm˜1, and a visualization of the prompt clusters in Figure˜9.
@@ 4 Experiments
This section presents a comprehensive empirical evaluation of $CRI$ compared to baseline initializations. We first present the experimental setting in Section˜4.1, and continue to discuss the results over the $HarmBench$ dataset in Section˜4.2. Additionally, we present ablation studies of our method in Section˜C.5, and extend the evaluations with additional attacks, models, and datasets in Section˜C.4. Our evaluation seeks to addresss three key research questions:
@@ 4.1 Experimental Setting
We present our experiments on the "standard behaviors" category of the $HarmBench$ dataset Mazeika et al. (2024), which is comprised of prompt-target pairs $(x,t)$. We define three disjoint subsets and utilize them in all the presented settings. $CRI$ utilizes $2$ subsets of $25$ samples each: a fine-tuning set for optimizing $CRI$, and a validation set for evaluating the universal attacks. The final subset is the $50$-sample test set over which we report the results. Additionally, we present results over the $AdvBench$ dataset Zou et al. (2023) in Sections˜C.4.3 and C.4.1.
Figure 4: Distribution of per-prompt Spearman correlation of LFSLFS and number of steps-to-success, on Llama-2. (Median $r=0.74$, $p=4.61\times 10^{-4}$).
@@
Figure 5: Comparison of ASRASR and number of steps on the $HarmBench$ dataset for the GCGGCG (top) and Embedding attacks (bottom) over the Llama-2 (left) and Vicuna (right) models.
@@
We evaluate our results over the open-source Vicuna-7B-v1.3 Chiang et al. (2023b), and Llama-2-7B-Chat-HF Touvron et al. (2023b) LLMs, which we accordingly denote as Vicuna, and Llama-2. We present these models to compare with the default setting presented by $GCG$, and provide results over additional models in Section˜C.1.
@@
We present the $GCG$ Zou et al. (2023) and the Embedding attacks Schwinn et al. (2024), where we utilized the same parameter configuration in all the presented settings, including the initializations’ training. The parameter configuration is then the default parameters suggested by the authors, except for the batch-size $B$ parameters of $GCG$. Thereby, $GCG$ utilizes $500$ optimization steps and the $JT$ set in Equation˜7 with $L=20$ tokens. Similarly, the Embedding attack utilizes $100$ optimization steps and the $JT$ set in Equation˜9 with $L=20$ tokens. To present a more robust setting, we limit the computational resources of the $GCG$ attack and take the batch-size parameter of $GCG$ to be $B=16$ instead of the default $512$ value. This entails that only $B=16$ tokens are sampled as possible replacements in each iteration of $GCG$, emphasizing the effect of the compared initializations. Additionally, we present results with increased batch size in Section˜C.5.1.
@@
We compare the KCRIK-CRI initializations over $K=1,5,25$ to existing approaches, i.e., "Standard" which utilizes repeated ’!!’ tokens, "Random" which utilizes uniform random tokens, and "Baseline" which utilizes randomly selected self-transfer initializations similarly to Andriushchenko et al. (2024); Jia et al. (2024). We then refer to "Standard" and "Random" initializations as uninformative and to others as informative. For each attack, $CRI$ pre-trains the transformations on the designated data subsets and utilizes $LFS$ to select an initialization for a given prompt. The "Baseline" approach then utilizes the same initializations set as $25\text{-}CRI$, but selects each initialization uniformly.
@@
In each attack’s evaluation, we follow the default evaluation framework suggested by the authors to compute the $ASR$. Thereby, $GCG$ evaluates success as not generating refusal-list keywords, where we provide the list in Section˜C.1.4. In contrast, the Embedding attack requires generating the exact target. We additionally present an evaluation of the resulting attacks via a GPT-4 judge in Section˜C.1.4. In addition to the $ASR$ and $LFS$, we define two metrics for estimating the computational overhead of attacks: MSSMSS—Median Steps to Success, ASSASS—Average Steps to Success. Hereby, a lower MSSMSS/ASSASS of initializations indicates their corresponding decrease in computational overhead.
@@
In all the presented settings, the computation was executed on Intel(R) Xeon(R) CPU and NVIDIA L40S GPU. An evaluation of runtime on this hardware can be found in Section˜C.5.2.